# Data-driven Virtual Test-bed of the Blown Powder Directed Energy Deposition Process

### v1

@@ 4 Conclusions
This study has presented a workflow which ingests OMM data to produce accurate sensor predictions for an unseen DED build. The model is able to produce quality predictions quickly, on the ms timescale. We have shown:
  • The predictive power of DMDc trained on OMM data
  • Produced control plots for operators and analysts using a relatively simple UQ
  • Demonstrated requirements for OMM recording
To further elaborate on the above points, the success of the DMDc representation relies on two critical factors: the density of time snapshots for adequate temporal resolution and the induced dynamics that ensure the collected data effectively capture the system’s behavior. Both elements must be carefully optimized to create an accurate and reliable model of the underlying physical process. Understanding the balance between interpolatory and extrapolatory behavior is essential in model construction, and despite the training and test datasets not achieving a uniform distribution, their similarity suggests that the model will primarily perform interpolation.
The state-space model was trained using the DMDc algorithm on approximately 1.5 million data points at a rate of $2.5\mu s$ per data point with an inference rate of $14.35\mu s$ per data point on relatively standard hardware (11th Gen Intel®i7 @ 3.00GHz). Results indicate robust model performance and these findings highlight the model’s reliability while identifying areas for further refinement.
The UQ results provide insights into noise characteristics and SNR trends for different observables in DED. While some trends align with expectations, others reveal unexpected challenges, particularly in working distance measurements. The proposed framework effectively combines model predictions with UQ to offer real-time bounds, assisting operators and analysts in maintaining process control and isolating anomalies. Although the noise frequency component of UQ does not directly enhance model predictions, it has potential utility in subsequent iterations of this model and workflow. Parity plot analysis provides additional context for areas of the models strengths and weaknesses, even suggesting portions of learned physics of the system.
Additionally, the study reveals that model accuracy in this workflow unsurprisingly degrades at lower data recording frequencies, with critical thresholds identified for various sensor observables. The analysis indicates that generally a minimum of at least one measurement is required spatially at the resolution of the laser spot diameter. These authors would like to emphasize the need for establishing minimum fidelity requirements for OMM in DED, shifting the focus from defect detection to ensuring sufficient sensor fidelity for accurate dynamical system modeling.
Finally, this workflow and results also represent an important first step towards feedback control of the observables in this study. In fact, a gain operator, $K_{gain}$, can easily be calculated from Equation 9 which casts the system in observable set points rather than raw process inputs, with the $y_{t}$’s coming from either previous experiments or simulated results.
$y_{t+1}$ $=\textbf{A}y_{t}+\textbf{B}u_{t}$ (12)
let $\text{let }u_{t}$ $=\textbf{K}_{gain}r_{t}\text{, where }r_{t}\text{is the referenceinput}$
$\textbf{K}_{gain}$ $=\frac{y_{t+1}-\textbf{A}y_{t}}{\textbf{B}r_{t}}$
@@ 5 Abbreviations
  • LENS®: Laser Engineered Net Shaping
  • LMD: Laser Metal Deposition
  • LSTM: Long, Short-Term Memory
  • LaSDI: Latent Space Dynamics Identification
  • OKID: Observer Kalman Filter Identification
@@
  • POD: Proper Orthogonal Decomposition
  • ROM: Reduced Order Modeling
  • SINDY: Sparse Identification of Nonlinear Dynamics
@@ 6 Acknowledgements
7 CRediT authorship contribution statement
Michael Juhasz: Writing - original draft, Conceptualization, Investigation, Methodology, Formal Analysis. Eric Chin: Writing - review & editing, Conceptualization. Youngsoo Choi: Writing - review & editing, Conceptualization. Joseph T. McKeown: Writing - review & editing, Resources, Funding acquisition, Project administration. Saad Khairallah: Writing - review & editing, Conceptualization, Methodology, Validation, Supervision.
@@ 8.1 Additional Figures
Figure S1: OMM data distributions for all other model inputs (features) in no particular order. Features are split into two separate plots and standardized for clarity. Shown are box plots comparing the training set to the test set with many features sharing distributions, however, some departures of the test set do exist.
Figure S2: DMDc results for training (a) and test data (b). (top) Time series of melt pool temperature for both ground truth and DMDc. Inset is the $R^{2}$ and mean squared error ($MSE$) values of the fit. (bottom) Residuals between the experimental data and model predictions.
Figure S3: DMDc results for training (a) and test data (b). (top) Time series of working distance for both ground truth and DMDc. Inset is the $R^{2}$ and mean squared error ($MSE$) values of the fit. (bottom) Residuals between the experimental data and model predictions.
Figure S4: Combined DMDc predictions with UQ applied as bounding envelope. Shown is working distance (mm). Notice the extent of the UQ envelope demonstrating an area of needed refinement.
8.2 The dataset
The raw dataset, detailed in section 2, comprises 26 DED experiments involving simple square prisms, part of a process parameter development study for Inconel 625. This raw dataset consisted of 49 columns/features and 1,706,677 rows/samples. Most features are listed in Table 1, with some omitted due to redundancy (e.g., Powder Hopper Disc Speed 1, 2, etc.). This raw dataset was examined for missing data (NaNs) and multi-collinearity (repetitive columns). No missing data was found, but would have removed by row/sample as part of a sound data analysis pipeline. Multi-collinearity was addressed using a feature selection algorithm based on VIF, reducing the dataset to 24 features. The test:train split for this study was 25 of those experiments for training and 1 for testing. Percentage-wise this roughly equates to a 9:91 test:train split. It is this final dataset which was used by both the DMDc algorithm and also for UQ. Notably, the dataset was not standardized before the split to avoid potential data bias.

### v2

@@ 4 Conclusions
This study has presented a workflow that effectively ingests OMM data to produce accurate sensor predictions for an unseen DED build. The model demonstrates the capability to generate quality predictions quickly, on the ms timescale. We have shown:
  • The predictive power of the DMDc trained, surrogate model using OMM data
  • Control plots for operators and analysts using a straightforward UQ approach
  • Key requirements for OMM data recording
To further elaborate on the above points, the success of the DMDc representation relies on two critical factors: the density of time snapshots for adequate temporal resolution and sufficiently exciting dynamics that ensure the collected data effectively capture the system’s behavior. Both elements must be optimized to create an accurate and reliable model of the underlying physical process.
The state-space, surrogate model was trained using the DMDc algorithm on approximately 1.5 million data points at a rate of $2.5\mu s$ per data point, achieving an inference rate of $14.35\mu s$ per data point on relatively standard hardware (11th Gen Intel®i7 @ 3.00GHz). Results indicate good model performance, highlighting the model’s reliability while identifying opportunities for further refinement.
The proposed framework effectively combines model predictions with UQ to offer realtime bounds, assisting operators and analysts in maintaining process control and isolating anomalies. Parity plot analysis provides additional context regarding the model’s strengths and areas for further exploration while also suggesting insights into the physics of the system.
Additionally, the study highlights that model accuracy in this workflow can vary with data recording frequencies, with critical thresholds identified for various sensor observables. The analysis underscores the importance of establishing minimum fidelity requirements for OMM in DED, shifting the focus to ensuring sufficient sensor fidelity for accurate dynamical system modeling.
@@ 5 Abbreviations
  • MLP: Multi-Layer Perceptron
@@ 7.1 Additional Figures
Figure S1: OMM data distributions for all other model inputs (features) in no particular order. Features are split into two separate plots and standardized for clarity. The complete list of final input features is as follows: Scan Rate (mm/min), Laser Commanded Power (Watts), Laser Power On Time (s), Heading Angle (deg), Part Energy Input (J), X Position (mm), Y Position (mm), Z Position (mm), C Position (deg), SLED ($J/mm^{2}$), Powder 2 Disc Speed (RPM), Contour Flag (0/1), NC Line Number (int), Enclosure Temperature (C), Enclosure H20 (ppm), Enclosure Pressure (mbar), Powder 2 Carrier Gas (l/min), Shield Gas (l/min), Current Slice (#), Program Slice Resolution (mm), and Post Layer Delay (s). Shown are box plots comparing the training set to the test set with many features sharing similar distributions, however, some departures of the test set do exist.
Figure S2: DMDc results for reproductive, or training, (a) and test data (b). (Left) Time series of melt pool temperature for both ground truth and DMDc state-space model. (Right) Histogram of residuals between the experimental data and model predictions.
Figure S3: DMDc results for reproductive, or training, (a) and test data (b). (Left) Time series of realtime working distance for both ground truth and DMDc state-space model. (Right) Histogram of residuals between the experimental data and model predictions.